Evaluate the integral of Csc² (2-3x) dx


1
Expert's answer
2021-03-09T02:13:00-0500

Solution:

"I=\int \csc^2 (2-3x)dx" ...(i)

Put "2-3x=t"

"\Rightarrow 0-3=\dfrac{dt}{dx}" [On differentiating]

"\Rightarrow dx=-\dfrac{dt}{3}"

Putting these values in (i), we get

"\Rightarrow I=\dfrac13\int(- \csc^2 t)dt"

"\Rightarrow I=\dfrac13\cot t+C"

"\Rightarrow I=\dfrac13\cot (2-3x)+C" [Substituting value of "t" back]